# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v3

@@ B.2 Proof of Lemma 2
Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$. For instance, if the $n$th ($n\in\{1,...,T-1\}$) order statistic of $z_{t}$ is $1.00$, its $n+1$th order statistic is $1.02$, and $M=2$, then all values of the threshold parameter $r_{1}$ between $1.00$ and $1.02$ correspond to the same model, as they all imply the same regime-switches. Hence, the identification of the threshold parameters should be understood in the sense that they are identified to be between some order statistics of the switching variable series $z_{t}$. Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. This concludes that the parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$. In other words, two distinct parameter sets give at most a measure-zero set of solutions for $(y_{1},...,y_{T})$.
@@
Logistic weights. Next, we show the ”moreover” part for the logistic transition weights defined in (2.4) with $M=2$. Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$. The proof proceed in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then we use these results to show that (3) the intercepts and autoregression matrices are uniquely identified.
@@
By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{t}=\phi_{t}$ and $\tilde{\boldsymbol{A}}_{t}=\boldsymbol{A}_{t}$ (irrespective of the weight function dynamics). By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, these equations can be written as
$\tilde{\phi}_{1,0}-\phi_{1,0}+\alpha_{2,t}(\phi_{1,0}-\phi_{2,0})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1,0}-\tilde{\phi}_{2,0})=0$ (B.7)
@@
Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in the (two) variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$. Therefore, for given coefficients, Equations (B.7) and (B.8) have at most one solution unless at least $d(1+pd)-1$ scalar equations are identically true. All the $d(1+pd)$ scalar equations are identically true only if $\phi_{1,0}=\phi_{2,0}$, $\tilde{\phi}_{1,0}=\tilde{\phi}_{2,0}$, $A_{1,i}=A_{2,i}$, and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$, which violates Condition (2) of Lemma 2. If exactly $d(1+pd)-1$ scalar equations are identically true, the coefficients for all but one of the scalar equations are zero, and $\alpha_{2,t}$ can be expressed in the form $\alpha_{2}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ for some constants $a_{0},a_{1}\in\mathbb{R}$. Since any single solution of $\alpha_{2,t},\tilde{\alpha}_{2,t}$ can be expressed in the form $\alpha_{2,t}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ for some $a_{0},a_{1}\in\mathbb{R}$, it follows that the solution to Equations (B.7) and (B.8) must satisfy $\alpha_{2,t}=a_{0}+a_{1}\tilde{\alpha}_{2,t}$ (for all $t$).
Making the dependence of the transition weights of $z_{t}$ explicit in the notation, observe that in the identity $\alpha_{2,t}(z_{t})=a_{0}+a_{1}\tilde{\alpha}_{2,t}(z_{t})$, the logistic functions $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are both continuous and strictly increasing maps of $z_{t}$ into $(0,1)$ (as $\gamma,\tilde{\gamma}>0$). Specifically, $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ increase smoothly from 00 to $1$ in an ”S”-shaped curve as $z_{t}$ goes from $\infty$- ∞ to $\infty$. Denote $h(z_{t})\equiv\alpha_{2,t}(z_{t})-a_{0}-a_{1}\tilde{\alpha}_{2,t}(z_{t})$, which is continuous in $z_{t}$ as $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are continuous.
Suppose $h(z_{t})=0$ is not identically true, and suppose $a_{1}\neq 1$. Then, $a_{1}\tilde{\alpha}_{2,t}(z_{t})$ takes values from 00 to $a_{1}$. If $a_{1}=0$, then $h(z_{t})=0$ at most in a single point by strict monotonicity of $\alpha_{2,t}(z_{t})$. If $a_{1}<0$, then $h(z_{t})$ is strictly increasing, so $h(z_{t})=0$ at most in a single point. If $a_{1}\in(0,1)$, then $a_{1}\tilde{\alpha}_{2,t}(z_{t})<\tilde{\alpha}_{2,t}(z_{t})$, whereas if $a_{1}>1$, then $a_{1}\tilde{\alpha}_{2,t}(z_{t})>\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}\in\mathbb{R}$. In either scenario, $h(z_{t})$ can cross zero only a finite number of times, as otherwise repeated crossing would force repeated sign changes, clearly contradicting the smooth logistic shapes that are “similar” but scaled. On the other hand, $h(z_{t})=0$ cannot hold on any set of positive Lebesgue measure because $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are strictly increasing logistic functions, and two such functions that coincide on an interval must coincide on $\mathbb{R}$, thereby forcing $a_{0}=0$ and $a_{1}=1$.
On the other hand, if $a_{1}=1$ and $a_{0}\neq 0$, then $\alpha_{2,t}(z_{t})$ and $a_{0}+a_{1}\tilde{\alpha}_{2,t}(z_{t})$ differ by constant shifting, and they can intersect in at most finitely many points. Consequently, $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ only in sets of Lebesgue measure zero in $z_{t}\in\mathbb{R}$, implying that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ (and thereby also $\alpha_{1,t}(z_{t})=\tilde{\alpha}_{1,t}(z_{t})$) almost everywhere in $z_{t}\in\mathbb{R}$. Thus, also $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ and $\alpha_{1,t}(z_{t})=\tilde{\alpha}_{1,t}(z_{t})$ for all $t$ almost everywhere in $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
@@
If $\tilde{\gamma}=\gamma$, it follows that $\tilde{c}=c$. If $\tilde{\gamma}\neq\gamma$, Equation (B.9) has only one solution, namely, $z_{t}=(\tilde{\gamma}-\gamma)^{-1}(\tilde{\gamma}\tilde{c}-\gamma c)$. Thus, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $z_{t}\in\mathbb{R}$.
We have established that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$, $\tilde{\gamma}=\gamma$, and $\tilde{c}=c$, almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$. Consequently, Equations (B.7) and (B.8) simplify to
$\tilde{\phi}_{1,0}-\phi_{1,0}+\alpha_{2,t}(\phi_{1,0}-\tilde{\phi}_{1,0})+\alpha_{2,t}(\tilde{\phi}_{2,0}-\phi_{2,0})=0$ (B.10)
@@
$\tilde{A}_{1,i}-A_{1,i}+\alpha_{2,t}(A_{1,i}-\tilde{A}_{1,i})+\alpha_{2,t}(\tilde{A}_{2,i}-A_{2,i})=0,\ \ i=1,...,p,$ (B.11)
respectively. Treating and $\phi_{m,0},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$ as fixed constants, Equations (B.10) and (B.11) define a system of $d(1+pd)$ linear equations and $2d(1+pd)$ variables ($\tilde{\phi}_{m,0},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,2$). Hence, there exists at most one solution to this system linear equations. Since one solution is $\tilde{\phi}_{m,0}=\phi_{m,0},\tilde{A}_{m,1}=A_{m,1},...,\tilde{A}_{m,p}=A_{m,p}$, $m=1,2$, this is the unique solution. This concludes the proof that the parameters $\phi_{m,0},A_{m,1},...A_{m,p}$, $m=1,...,M$ as well as the parameter $r_{1},...,r_{m-1}$ models with threshold and the parameters $c,\gamma$ for models with logistics transition weights are uniquely identified almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$ and $z_{t}\in\mathbb{R}^{T}$.
Exogenous nonrandom weights. Suppose the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are exogenous (nonrandom). Then, arguments similar to those used for logistic weights, but applied for any finite number of regimes, show that the parameters $\phi_{m,0},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in almost everywhere in $(y_{1},...,y_{T})\in\mathbb{R}^{dT}$.∎

### v4

@@ B.2 Proof of Lemma 2
Because a regime switches when (and only when) the switching variable passes one of the threshold parameter values, for a given (countable number of) observations of $z_{t}$, the threshold parameters can only be identified with the resolution determined by the spacing of the order statistics of $z_{t}$. For instance, if the $n$th ($n\in\{1,...,T-1\}$) order statistic of $z_{t}$ is $1.00$, its $n+1$th order statistic is $1.02$, and $M=2$, then all values of the threshold parameter $r_{1}$ between $1.00$ and $1.02$ correspond to the same model, as they all imply the same regime-switches. Hence, the identification of the threshold parameters should be understood in the sense that they are identified to be between some order statistics of the switching variable series $z_{t}$. Now, if the threshold parameter values $\tilde{r}_{1},...,\tilde{r}_{M-1}$ are between different order statistics of $z_{t}$ than $r_{1},...,r_{M-1}$, the partitions $\mathcal{\tilde{T}}_{m}$ and $\mathcal{T}_{m}$ are different for some $m\in\{1,...,M\}$, in which case, by the above discussion, Equation (B.6) can only hold in sets of Lebesgue measure zero in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$. This concludes that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$ and $r_{1},...,r_{M-1}$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
@@
Logistic weights. Next, we show the ”moreover” part for the logistic transition weights defined in (2.4) with $M=2$. Suppose Equation (B.4) holds for the two sets of parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,...,M$, $c,\gamma$ and $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,...,M$, $\tilde{c},\tilde{\gamma}$. The proof proceeds in the following steps: (1) we show that $\tilde{\alpha}_{m,t}=\alpha_{m,t}$ for each $t$ almost everywhere in $z_{t}\in\mathbb{R}$, (2) we use this identity to show that the parameters $c,\gamma$ are uniquely identified (almost everywhere), and then (3) we use these results to show that the intercepts and autoregression matrices are uniquely identified.
@@
By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$ (irrespective of the weight function dynamics). By making use of the identity $\alpha_{1,t}=1-\alpha_{2,t}$, these equations can be written as
$\tilde{\phi}_{1}-\phi_{1}+\alpha_{2,t}(\phi_{1}-\phi_{2})-\tilde{\alpha}_{2,t}(\tilde{\phi}_{1}-\tilde{\phi}_{2})=0$ (B.7)
@@
Equations (B.7) and (B.8) define a system of $d(1+pd)$, $d\geq 2,p\geq 1$, linear scalar equations in two variables $\alpha_{2,t}$ and $\tilde{\alpha}_{2,t}$ (interpreting the AR parameters as coefficients here). Suppose first that all of them are identically true, i.e., that all the coefficients are zero. In (B.7), this requires that $\tilde{\phi}_{1}=\phi_{1}$, $\phi_{1}=\phi_{2}$, and $\tilde{\phi}_{1}=\tilde{\phi}_{2}$. Similarly in (B.8), zero coefficients require that $\tilde{A}_{1,i}=A_{1,i}$, $A_{1,i}=A_{2,i}$ and $\tilde{A}_{1,i}=\tilde{A}_{2,i}$ for all $i=1,...,p$. But as $\phi_{1}=\phi_{2}$ and $A_{1,i}=A_{2,i}$ for all $i=1,...,p$ violates Condition (2) of Lemma 2, it must be that not all scalar equations in (B.7) and (B.8) are identically true.
@@
For each scalar equation that is not identically true, the coefficients are not all zero, and each of these scalar equations can be expressed in the form $a_{2,j}\alpha_{2,t}=a_{0,j}+a_{3,j}\tilde{\alpha}_{2,t}$ for some $a_{0,j},a_{2,j},a_{3,j}\in\mathbb{R}$ for the $j$th not identically true equation. If $a_{2,j}=0$, then $\tilde{\alpha}_{2,t}$ is the constant $-a_{0,j}/a_{3,j}$ for all $t$, which violates Condition (3) of Lemma 2. Hence, $a_{2,j}\neq 0$ (in the not identically true equations), implying that $\alpha_{2,t}=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}$, where $a_{1,j}\equiv a_{3,j}a_{2,j}^{-1}$. Thus, the solution to Equations (B.7) and (B.8) must satisfy this relationship (for all $j$).
Making the dependence of the transition weights of $z_{t}$ explicit in the notation, suppose $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds in a set of positive Lebesque measure in $z_{t}\in\mathbb{R}$. Since $\alpha_{2,t}(z_{t})$ and $\tilde{\alpha}_{2,t}(z_{t})$ are logistic, $\alpha_{2,t}(z_{t})$ and $a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ are real analytic functions. Thus, by the Identity theorem, we have that $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ for all $z_{t}$. Since this must hold for all $z_{t}$, it follows from $\underset{z_{t}\rightarrow-\infty}{\lim}\alpha_{2,t}(z_{t})=0$ and $\underset{z_{t}\rightarrow-\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}$ that $a_{0,j}=0$. Similarly, assuming $a_{1,j}\neq 0$, it follows from $a_{0,j}=0$, $\underset{z_{t}\rightarrow+\infty}{\lim}\alpha_{2,t}(z_{t})=1$, and $\underset{z_{t}\rightarrow+\infty}{\lim}(a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t}))=a_{0,j}+a_{1,j}$ that $a_{1,j}=1$. If $a_{1,j}=0$, then clearly $\alpha_{2,t}(z_{t})=a_{0,j}+a_{1,j}\tilde{\alpha}_{2,t}(z_{t})$ holds only in the limit $z_{t}\rightarrow-\infty$, and thereby not in a set of positive Lebesgue measure. Thus, it must be that $a_{0,j}=0$ and $a_{1,j}=1$ in sets of positive Lebesgue measure, concluding that $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ almost everywhere in $z_{t}\in\mathbb{R}$.
@@
If $\tilde{\gamma}=\gamma$, it follows that $\tilde{c}=c$. If $\tilde{\gamma}\neq\gamma$, Equation (B.9) has only one solution, namely, $z_{t}=(\tilde{\gamma}-\gamma)^{-1}(\tilde{\gamma}\tilde{c}-\gamma c)$. Thus, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ and $z_{t}\in\mathbb{R}$.
We have shown that for each $t$, $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ and $z_{t}\in\mathbb{R}$. If the $z_{t}$ is a function of $\boldsymbol{y}_{t-1}$ (rather than exogenous), we have that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ (as a countable union of measure zero sets has a measure zero). Since this holds for each $t$, combining it for $t=1,....,T$ yields that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. If $z_{t}$ is a function of $\boldsymbol{y}_{t-1}$, this result holds almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
Before showing that the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, are identified, we present the following lemma that is used in the proof:
Lemma B.2.
Consider the $T$ linear equations in $M$ variables
$\alpha_{1,t}x_{1}+...+\alpha_{M,t}x_{M}=c_{t},\quad t=1,...,T,$ (B.10)
where $x_{1},...,x_{M}$ are scalar variables, and $\alpha_{1,t},...,\alpha_{M,t}$ and $c_{t}$ are scalar constants. Suppose there exists $M$ indices $t_{1},...,t_{M}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent. Then, there exists at most one solution to the system of linear equations (B.10).
Proof.
By assumption, there exists $M$ indices $t_{1},...,t_{M}$ such that the coefficient vectors $(\alpha_{1,t_{i}},...,\alpha_{M,t_{i}})$, $i=1,...,M$, are linearly independent. Now, the rows of the $(M\times M)$ matrix
$A^{\alpha}\equiv\begin{bmatrix}\alpha_{1,t_{1}}&\cdots&\alpha_{M,t_{1}}\\ \vdots&\ddots&\vdots\\ \alpha_{1,t_{M}}&\cdots&\alpha_{M,t_{M}}\\ \end{bmatrix}$ (B.11)
are linearly independent, and thus $A^{\alpha}$ is invertible. The $M$ linear equations corresponding to these indices form the system $A^{\alpha}x=c$, where $x=(x_{1},...,x_{M})$ and $c=(c_{t_{1}},...,c_{t_{M}})$. Because $A^{\alpha}$ is invertible, this system has exactly one solution.
Now, any solution $x$ of the full system (B.10) for $t=1,....,T$ must satisfy the above subsystem, which has exactly one solution. Therefore, the full system (B.10) has at most one solution. ∎
We have established that $\tilde{\gamma}=\gamma$ and $\tilde{c}=c$ (and $\alpha_{2,t}(z_{t})=\tilde{\alpha}_{2,t}(z_{t})$ for all $t$) almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$. Using these identities, Equations (B.7) and (B.8) simplify to
$(1-\alpha_{2,t})\tilde{\phi}_{1}+\alpha_{2,t}\tilde{\phi}_{2}=(1-\alpha_{2,t})\phi_{1}+\alpha_{2,t}\phi_{2}$ (B.12)
@@
$(1-\alpha_{2,t})\tilde{A}_{1,i}+\alpha_{2,t}\tilde{A}_{2,i}=(1-\alpha_{2,t})A_{1,i}+\alpha_{2,t}A_{2,i},\ \ i=1,...,p,$ (B.13)
respectively. Treating the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$, as well as ($\mathcal{F}_{t-1}$-measurable or nonrandom) transition weights $\alpha_{2,t}$ as coefficients here, at each $t$, each scalar equation in (B.12)-(B.13) is a linear equation in two variables given by the corresponding entries of $\tilde{\phi}_{m},\tilde{A}_{m,1},...,\tilde{A}_{m,p}$, $m=1,2$.
Each scalar equation in (B.12)-(B.13) defined together for all $t=1,...,T$ defines a system of $T$ linear equations and two variables. By Condition (3) of Lemma 2 (with $M=2$), there exists two indices $t_{1},t_{2}\in\{1,....,T\}$ such that the corresponding coefficient vectors $(\alpha_{1,t_{i}},\alpha_{2,t_{i}})$, or equally $(1-\alpha_{2,t_{i}},\alpha_{2,t_{i}})$, $i=1,2$, are linearly independent. Hence, it follows from Lemma B.2 that there exists at most one solution to each such system of scalar equations, and thereby also to the system of $Td(1+pd)$ linear equations (B.12)-(B.13), $t=1,....,T$, in $2d(1+pd)$ variables. Since one solution is $\tilde{\phi}_{m}=\phi_{m},\tilde{A}_{m,i}=A_{m,i}$, $i=1,...,p$, $m=1,2$, this is the unique solution. This concludes the proof that the parameters $\phi_{m},A_{m,1},...A_{m,p}$, $m=1,...,M$ and $c,\gamma$ are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$ and, if the switching variable $z_{t}$ is exogenous, $(z_{1},...,z_{T})\in\mathbb{R}^{T}$.
Exogenous nonrandom weights. Suppose the transition weights $\alpha_{1,t},...,\alpha_{M,t}$ are exogenous (nonrandom). If the transition weights are binary, i.e., $\alpha_{m,t}\in\{0,1\}$ for all $m=1,...,M$ and $t=1,...,T$, then arguments similar to those used for threshold weights can be used to conclude that the parameters $\phi_{m},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.
Next, suppose the transition weights are not binary. By applying Lemma B.1 to each scalar equation in Equation (B.4), it follows that Equation (B.4) holds almost everywhere in $\boldsymbol{y}_{t-1}\in\mathbb{R}^{dp}$ if only if Equation (B.4) is identically true, that is, $\tilde{\phi}_{y,t}=\phi_{y,t}$ and $\tilde{\boldsymbol{A}}_{y,t}=\boldsymbol{A}_{y,t}$. Substituting the exogenous weights ($\alpha_{m,t}=\tilde{\alpha}_{m,t}$ is given for all $m=1,...,M$ and $t=1,...,T$) to these equations yields
$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{\phi}_{m}-\phi_{m})$ $=0,\quad\text{and}$ (B.14)
$\sum_{m=1}^{M}\alpha_{m,t}(\tilde{A}_{m,i}-A_{m,i})$ $=0,\quad i=1,...,p,$ (B.15)
respectively.
Treating the parameters $\phi_{m},A_{m,1},...,A_{m,p}$, $m=1,2$, and $c,\gamma$, as well as transition weights $\alpha_{1,t},...,\alpha_{M,t}$ as constants, each scalar equations in (B.14)-(B.15) defined together for all $t=1,...,T$ defines a system of $T$ linear equations in $M$ variables. Therefore, arguments similar to those used for logistic weights but applied for any finite number of regimes regimes, making use of Lemma B.2, can be used to conclude that the parameters $\phi_{m},A_{m,1},...,A_{M,p}$, $m=1,...,M$, are uniquely identified almost everywhere in $(y_{-p+1},...,y_{0},y_{1},...,y_{T-1})\in\mathbb{R}^{d(T+p-1)}$.∎